A gear 12 cm in diameter, is turning a gear 18 cm in diameter. when the smaller gear has made 42 revolutions how many has the larger one made?
a)24
b)28
c)50
d)63
e)72

Answer by stanbon(75887)

A gear 12 cm in diameter, is turning a gear 18 cm in diameter. when the smaller gear has made 42 revolutions how many has the larger one made?
----
smaller gear perimeter: 12(pi) in
larger gear perimeter: 18(pi) in
----
# of revs of larger = (42*12(pi))/18(pi) = (2/3)42 = 28
==============
Cheers,
Stan H.
==============
a)24
b)28
c)50
d)63
e)72